OpenStudy (anonymous):

a shell of mass o.020 kg is fired by gun of 100kg the muzzle speed of the shell is 10 m/s what is the recoil of speed of gun

OpenStudy (anonymous):

by the law of conservation of momentum, we have m1*v1=m2*v2 or, 0.020*10=100*v2 or, v2=0.002m/s Thus, the recoil speed of gun is 0.002m/s.